# Conformal Predictive Systems Under Covariate Shift

### v1

@@ Appendix A Split Conformal Predictive System
For Split CPS (SCPS), the same procedure is followed as a split conformal prediction; the training sequence $z_{1:n}$ is split into two: a proper training sequence $z_{1:m}$ and calibration sequence $z_{m+1:n}$. Similarly as an CPS, an SCPS is defined as a function that is both a split conformal transducer (Definition 4) and an RPS (Definition 2) [Vovk et al., 2020a].
Definition 3 (Inductive (Split) Conformity Measure, Vovk et al. [2022]).
@@
Definition 4 (Split Conformal Transducer, Vovk et al. [2020a]).
The split conformal transducer determined by a split conformity measure $A$ (see Definition 3) is defined as,
$\begin{split}Q(z_{1},...,z_{n},(x,y),\tau):=&\sum_{i=m+1}^{n}[R_{i}<R^{y}]\frac{1}{n-m+1}\\ &+\sum_{i=m+1}^{n}[R_{i}=R^{y}]\frac{\tau}{n-m+1}\\ &+\frac{\tau}{n-m+1}\end{split}$ (19)
@@
$\begin{split}R_{i}&:=A(z_{1},...,z_{m},(x_{i},y_{i})),\qquad i=m+1,...,n,\\ R^{y}&:=A(z_{1},...,z_{m},(x,y)),\qquad y\in\mathbb{R}.\end{split}$
Vovk et al. [2020a] proofs that any split conformal transducer is an RPS if and only if it is based on a balanced isotonic split conformity measure (Definition 6).
Definition 5 (Isotonic Split Conformity Measure, Vovk et al. [2020a]).
@@
Definition 6 (Balanced Isotonic Split Conformity Measure, Vovk et al. [2020a]).
An isotonic split conformity measure $A$ (see Definition 5) is balanced if, for any $m$ and $z_{1},...,z_{m}$, the set
@@ Appendix B Python Package: crepes-weighted
For the simulation experiments in this work, we implemented the proposed WSCPS and the WCP [Tibshirani et al., 2019] in crepes-weighted, which is an extension of crepes [Boström, 2022], a Python package that implements conformal classifiers, regressors, and predictive systems on top of any standard classifier and regressor. crepes-weighted relies on the same classes and functions as crepes, with the slight modification that for the ConformalRegressor and ConformalPredictiveSystem classes, the methods fit and predict needs to include the likelihood ratios of each calibration and test object respectively.

### v2

@@ Appendix A Split Conformal Predictive System
For split CPS (SCPS), the same procedure is followed as a split conformal prediction; the training sequence $z_{1:n}$ is split into two: a proper training sequence $z_{1:m}$ and calibration sequence $z_{m+1:n}$. Similarly as an CPS, an SCPS is defined as a function that is both a split conformal transducer (Definition A.2) and an RPS (Definition 2.2) (Vovk et al., 2020a).
Definition A.1 (Inductive (Split) Conformity Measure, Vovk et al. (2022)).
@@
Definition A.2 (Split Conformal Transducer, Vovk et al. (2020a)).
The split conformal transducer determined by a split conformity measure $A$ (see Definition A.1) is defined as,
$\begin{split}Q(z_{1},...,z_{n},(x_{n+1},y),\tau):=&\sum_{i=m+1}^{n}[R_{i}<R^{y}_{n+1}]\frac{1}{n-m+1}\\ &+\sum_{i=m+1}^{n}[R_{i}=R^{y}_{n+1}]\frac{\tau}{n-m+1}\\ &+\frac{\tau}{n-m+1}\end{split}$ (19)
@@
$\begin{split}R_{i}&:=A(z_{1},...,z_{m},(x_{i},y_{i})),\qquad i=m+1,...,n,\\ R^{y}_{n+1}&:=A(z_{1},...,z_{m},(x_{n+1},y_{n+1})),\qquad y\in\mathbb{R}.\end{split}$
Vovk et al. (2020a) proofs that any split conformal transducer is an RPS if and only if it is based on a balanced isotonic split conformity measure (Definition A.4).
Definition A.3 (Isotonic Split Conformity Measure, Vovk et al. (2020a)).
@@
Definition A.4 (Balanced Isotonic Split Conformity Measure, Vovk et al. (2020a)).
An isotonic split conformity measure $A$ (see Definition A.3) is balanced if, for any $m$ and $z_{1},...,z_{m}$, the set
@@ Appendix B Python Package: crepes-weighted
For the simulation experiments in this work, we implemented the proposed WSCPS and the WCP (Tibshirani et al., 2019) in crepes-weighted, which is an extension of crepes (Boström, 2022), a Python package that implements conformal classifiers, regressors, and predictive systems on top of any standard classifier and regressor. crepes-weighted relies on the same classes and functions as crepes, with the slight modification that for the ConformalRegressor and ConformalPredictiveSystem classes, the methods fit and predict needs to include the likelihood ratios of each calibration and test object respectively.